Given the frequency table below, what is the probability that a respondent is a child given that the respondent prefers Brand 1? (keep two decimals)

brand 1 brand 2
children 25 75 100
adults 30 70 100
total 55 145 200
Answer #1

Probably of those who prefer brand 1

= 55/200

Probability of a child preferring brand 1 is

=25/55

combining these two conditions together, the answer becomes

=(55/200) * (25*55)

= 25/200

= 0.125